two boats leave san diego at the same time. one boat sails due north and the due west. when they are 10 miles apart, the boat traveling north has sailed 2 miles farther than the boat sailing west.find the distance each boat has traveled.
Answer by mananth(16946)

The boats are travelling at right angles
At any given time if you draw a line from the starting point to the boats and draw a line joining the two boats they form a right triangle
distance = 10 miles this is the hypotenuse
let west bound boat be at x miles from starting point
other boat is x+2 miles from origin
x^2+(x+2)^2=10^2
x^2+x^2+4x+4=100
2x^2+4x-96 = 0
/2
x^2+2x-48=0
x^2+8x-6x-48=0
x(x+8)-6(x+80=0
(x+8)(x-6)=0
S0 x=6
west bound boat is 6 miles
North bound boat is 8 miles from origin